OpenStudy (anonymous):

How many solutions does the system have? x=-4y+4 2x+8y=8

OpenStudy (anonymous):

-4 = - 4y -x --> I) 4 = 4y +x II) 8 = 8y + 2x we multiply the I) with 2 8 = 8y + 2x 8 = 8y + 2x and subtract them both 0 = 0

jhonyy9 (jhonyy9):

answer ?

OpenStudy (anonymous):

one solution, the equations are the same. divide the second one by 2 to know.

OpenStudy (anonymous):

I thought it was 2 solutions